OpenStudy (anonymous):

What is the solution to 2 square root of x-8=6?

OpenStudy (anonymous):

\[2\sqrt{x-8}=6\] \[\sqrt{x-8}=3\] \[x-8=3^2=9\] one more step

OpenStudy (anonymous):

is the answer 1?

hero (hero):

no

OpenStudy (anonymous):

so whats the answer??? im confused

hero (hero):

about what?

OpenStudy (anonymous):

i got 1 from subtracting 8 in the equation x-8=3^2

OpenStudy (anonymous):

first i squared the 3 and got 9 then i subtracted 8 and got 1

hero (hero):

Aren't you supposed to add eight to both sides?

OpenStudy (anonymous):

19+8=17 is the ans

OpenStudy (anonymous):

oh! i completely missed that! thankyou

hero (hero):

yeah, thanks Raja but I was clearly asking ashley

hero (hero):

By the way, \(19 + 8 \ne 17\)

hero (hero):

@Raja99

OpenStudy (anonymous):

sry..9+8=17